If the sides of a square are lengthened by 7 cm, the area becomes 121 cm^2. find the length of a side of the original square
Answer by jorel1380(3719)

(x+7)(x+7)=121
x2+14x+49=121
x2+14x-72=0
(x+18)(x-4)=0
Throwing away the negative answer, we get x=4..